OpenStudy (anonymous):

The number of subscribers y to a website after t years is shown by the equation below: y = 50(1.75)^t Which conclusion is correct about the number of subscribers to the website?

OpenStudy (anonymous):

(A) It increased by 75% every year. (B) It decreased by 75% every year. (C) It increased by 50% every year. (D) It decreased by 50% every year.

OpenStudy (anonymous):

Is this A? @ganeshie8

OpenStudy (anonymous):

@hartnn @iambatman @pooja195

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

YW, Also this equation is a growth and decay equation .

OpenStudy (anonymous):

Also, in the equation you can tell it is going to be 75% because this: y = 50(1.75)^t = y = 50(1 + .75)^t which .75 is 75 % and it grows and not decays